A person walks 750m due north, then 250m due east. If the entire walk


takes 13 minutes, find the person's average velocity?



An object moves along a straight line. First it travels at a velocity of


12m/s for 6s and then continues at the same direction with 20m/s for 3s.


What is its average speed?



For 10s, the velocity of a car which travels with a constant acceleration,


changes from 10m/s to 20m/s. How far does the car travel?



A cyclist covers a distance of 15 miles in 2 hours. Calculate his speed.



A boy walks at a speed of 4 kmph. How much time does he take to walk a


distance of 20 km?



A car takes 4 hours to cover a distance, if it travels at a speed of 40 mph. What should be its speed to cover the same distance in 1.5 hours?



If a person drives his car in the speed 50 miles per hour, how far can he


cover in 90 minutes?

1
Expert's answer
2022-04-27T16:32:29-0400

a.


"v_{ave}=\dfrac{750m+250m}{13min}=\dfrac{1000}{13}m/min"

b.


"v_{ave}=\dfrac{12m/s(6s)+20m/s(3s)}{6s+3s}=\dfrac{44}{3}m/s"

c.


"v_1=v_0+at_1=>a=\dfrac{v_1-v_0}{t_1}"

"s_1=v_0t_1+\dfrac{at_1^2}{2}"

"s_1=v_0t_1+\dfrac{(v_1-v_0)t_1}{2}"

"s_1=\dfrac{(v_1+v_0)t_1}{2}"

"s_1=\dfrac{(20m/s+10m/s)10s}{2}=150m"

d.


"v=15mi/2h=7.5\ mi/h"

e.


"t=\dfrac{20km}{4km/h}=5hours"

f.


"v=\dfrac{40mi/h(4h)}{1.5h}=\dfrac{320}{3}mi/h"

"\dfrac{320}{3}" mph


g. "90min=1.5h"


"50mi\/h(1.5h)=75mi"

75 miles